A buffer solution contains 0.330 M ammonium chloride and 0.319 M ammonia.

If 0.0605 moles of potassium hydroxide are added to 250 mL of this buffer, what is the pH of the resulting solution ?
(Assume that the volume does not change upon adding potassium hydroxide)

pH =
1
Expert's answer
2020-10-30T06:42:43-0400

pOH = pKb + log([NH4+]/[NH3])

pOH = -log(1.8e-5) + log(0.330/0.319)

pOH = 4.76 - before adding NaOH.

Added NaOH: 0.0605 / 0,25 = 0.242 M - concentration

pOH = pKb + log(0.330-0.242/0.242+0.319)

pOH = 3.94 - after adding NaOH

pH = 14 - pOH = 14 - 3.94 = 10.06